The amplitude of a wave at a distance of 5.0 m from a small point source is 0.20 mm. What would be

the amplitude if the distance from the source is increased to 20 m?


1
Expert's answer
2021-10-24T18:23:52-0400

Given:

"r_1=5.0\:\rm m"

"r_2=20\:\rm m"

"A_1=0.20\:\rm mm"


The wave from a point source has spherical wavefront, so its amplitude is inversely proportional to the distance. Hence


"\frac{A_1}{A_2}=\frac{r_2}{r_1}"

"A_2=A_1\frac{r_1}{r_2}=\rm0.20\: mm*\frac{5.0\: m}{20\: m}=0.05\: mm"